OpenStudy (anonymous):

factor by grouping. 12x^2-9xy-8xy+6y^2

OpenStudy (mathmale):

Julie: How have you applied factoring by grouping in the past? Note how the first two terms both contain x and that the last two terms both contain y. This suggests that x can be factored out of the first two terms and y out of the last two terms: x(12x - 9y) - y(8x - 6y). Now look at that (12x - 9y) and (8x - 6y). See any pattern? See anything in common to both? Note how (12x - 9y) = 3(4x - 3y) and (8x - 6y) = 2(4x - 3y). See anything / any factor common to both? Hope this helps you get started at factoring 12x^2-9xy-8xy+6y^2 by grouping. Please give it a try on your own. Further questions? Just ask.

OpenStudy (anonymous):

is (12x^2+6y^2)(-9xy-8xy) right

OpenStudy (mathmale):

You could check your answer by multiplying your two factors together.

OpenStudy (mathmale):

Let's go back to x(12x - 9y) - y(8x - 6y). Inside the first set of parentheses is 12x - 9y. Do these terms have anything in common? If there's a common factor, please factor it out. Inside the second set of parentheses is 8x - 6y. Do these terms have anything in common? If there's a common factor, please factor it out.

OpenStudy (anonymous):

3x-3y & 2y-3?

OpenStudy (mathmale):

Let's go back to x(12x - 9y) - y(8x - 6y). Would you agree that, after factoring, this expression becomes 3x(4x - 3y) - 2y(4x - 3y)? Again, I ask you to look for any term or factor that shows up more than once and thus is a common factor. After factoring, 3x(4x - 3y) - 2y(4x - 3y) becomes (4x - 3y)*( what? )

OpenStudy (anonymous):

12x^2-9xy

OpenStudy (mathmale):

Julie, it appears that you are MULTIPLYING where you should be doing the opposite: factoring. You've multiplied 3x*(4x - 3y). Instead, please look at 3x(4x - 3y) - 2y(4x - 3y) again and verify the following: (4x - 3y) is the factor common to both the left and the right term of this expression. Factoring (4x - 3y) out, we are left with (3x - 2y). Thus, 3x(4x - 3y) - 2y(4x - 3y) factors into (4x - 3y)(3x - 2y). Check by multiplying this out.

OpenStudy (anonymous):

thanks

OpenStudy (mathmale):

You're welcome! Hope you can successfully apply these principles to other factoring by grouping problems.